# mDPO: Conditional Preference Optimization for Multimodal Large Language Models

### v1

@@ 1 Introduction
Direct preference optimization (DPO) Rafailov et al. (2023) has emerged as the predominating method for aligning large language models (LLMs) with human preferences Rafailov et al. (2023). Building on its success in the language modality, recent studies have extended DPO to multimodal scenarios Li et al. (2023); Yu et al. (2024a); Zhou et al. (2024); Zhao et al. (2023). However, transferring this approach across modalities presents significant challenges. Merely substituting textual preference data with multimodal preference data does not consistently yield positive outcomes and can exacerbate issues such as hallucinations Li et al. (2023); Sarkar et al. (2024).
@@
While recent efforts in multimodal preference learning focus on improving performance through enhanced multimodal preference data Li et al. (2023); Zhao et al. (2023); Xiao et al. (2024); Zhou et al. (2024); Pi et al. (2024); Sarkar et al. (2024); Yu et al. (2024b); Deng et al. (2024), we investigate the pitfalls of multimodal DPO from a different perspective. Through controlled comparisons, we discover that multimodal LLMs can achieve similar performance even when all images are removed from the multimodal preference data during DPO (see Figure 1). This counterintuitive finding suggests that the failure of DPO in multimodal scenarios may not be solely attributed to data quality. We attribute this to a systematic gap between the theoretical expectations and practical implementations of the DPO objective in multimodal settings (refer to Figure 2). While DPO aims to compute implicit rewards conditioned on all input modalities, it may prioritize language-only preferences and overlook the image condition (i.e., unconditional preference), leading to suboptimal model performance and increased hallucination. After applying DPO, the model may show an increased tendency to ignore the provided image and generate responses based solely on the question (illustrated in Figure 3).
@@ 3.1 Conditional Preference Optimization
$\mathcal{L}_{\text{CoPO}}=-\log\sigma\Big{(}\beta\log\frac{\pi{\theta}(y_{w}|m_{w},q)}{\pi_{\text{ref}}(y_{w}|m_{w},q)}-\beta\log\frac{\pi_{\theta}(y_{w}|m_{l},q)}{\pi_{\text{ref}}(y_{w}|m_{l},q)}\Big{)}.$
@@ 4.1 Experimental Setup
Models. We apply mDPO on two multimodal LLMs in different sizes. Bunny-v1.0-3B He et al. (2024) is a 3B model building upon SigLIP Zhai et al. (2023) and Phi-2 Javaheripi et al. (2023). It is pretrained on 2M image-text pairs and finetuned on 1.3M instruction tuning data. LLAVA-v1.5-7B Liu et al. (2024a) is a 7B model based on CLIP Radford et al. (2021) and Vincuna Chiang et al. (2023). It is pretrained on 558K image-text pairs and finetuned on 665K instruction tuning data.
Preference Data. We sample 10K preference data from Silkie Li et al. (2023) with instructions from LLaVA-Instruct-150K Liu et al. (2024a) for training. The origina Silkie dataset contains 80K preference data collected on 12 multimodal LLMs. While the original Silkie paper explores the effect of extreme data size, we follow the majority of prior works using around 10K data for preference optimization Sun et al. (2023); Zhao et al. (2023).
@@
Baselines. We primarily compare mDPO with standard DPO. The standard DPO baseline shares the same training process, data, and hyper-parameters, despite different learning objectives. We further provide the results of other multimodal LLMs for reference, although they are not directly comparable due to different base models, preference data, and alignment methods. This group contains GPT-4V Achiam et al. (2023), LLaVA-v1.5-13B Liu et al. (2024a), Qwen-VL-Chat Bai et al. (2023), POVID Zhou et al. (2024), HACL Jiang et al. (2024), OPERA Huang et al. (2024), VCD Leng et al. (2024), EOS Yue et al. (2024), HA-DPO Zhao et al. (2023), HALVA Sarkar et al. (2024), RLHF-V Yu et al. (2024a), HSA-DPO Xiao et al. (2024), and Silkie Li et al. (2023). Some of them are contemporary works to ours.
@@ 5 Related Work
Reinforcement learning from human feedback (RLHF; Christiano et al. 2017; Ouyang et al. 2022) has proven to be an effective approach for aligning LLMs with human values. Direct preference optimization (DPO; Rafailov et al. 2023), which involves directly optimizing LLMs based on human preferences, has been widely adopted in RLHF due to its strong performance and the elimination of the need for a separate reward model. Significant efforts have been made to further enhance the efficacy and efficiency of DPO, which can be categorized into algorithmic and data-related advancements. On the algorithmic side, various approaches aim to improve the efficiency of DPO. For example, ORPO Hong et al. (2024) models preferences using an odds ratio and combines instruction fine-tuning and preference optimization into a unified training process. Methods such as CPO Xu et al. (2024), TPO Saeidi et al. (2024), and SimPO Meng et al. (2024) simplify DPO by eliminating the use of a reference model, thereby reducing computational and memory overhead. Additionally, IPO Azar et al. (2024) addresses the issue of reward overfitting in DPO. On the data side, approaches such as KTO Ethayarajh et al. (2024) and NCA Chen et al. (2024) seek to overcome DPO’s requirement for paired preference data by designing optimization goals that can also utilize unpaired data. Iterative DPO Xu et al. (2023); Yuan et al. (2024) and SPPO Wu et al. (2024) propose sampling preference data in an on-policy manner, achieving better results than off-policy DPO. More closely related to our work, studies by Park et al. (2024) and Dong et al. (2024) address the reward hacking problem in textual preference optimization, where human preference may be biased towards longer outputs. They propose to calibrate the rewards in DPO with respect to output length. In this work, we discover a novel challenge in multimodal DPO, where preference optimization often neglects images. We then propose a solution to this problem through conditional preference optimization.
In multimodal scenarios, recent works mainly focus on creating multimodal preference data Li et al. (2023); Zhao et al. (2023); Xiao et al. (2024); Zhou et al. (2024); Pi et al. (2024); Sarkar et al. (2024); Yu et al. (2024b); Deng et al. (2024). These efforts include collecting human preference Sun et al. (2023); Yu et al. (2024a), preference from advanced multimodal LLMs Li et al. (2023); Yu et al. (2024b), and preference from the model to align itself Deng et al. (2024). In terms of learning objectives, recent works mainly follows DPO for LLMs Li et al. (2023); Zhao et al. (2023); Zhou et al. (2024). Some also apply reinforcement learning Sun et al. (2023) and contrastive learning Sarkar et al. (2024); Jiang et al. (2024). Our work studies an overlooked but crucial problem in the multimodal DPO objective.
@@ 6 Conclusion
We propose mDPO, a preference optimization method dedicated to multimodal scenarios. mDPO leverages conditional preference optimization to encourage multimodal LLMs to capture preference labels based on both visual and language cues. It further introduces anchored preference optimization to prevent the likelihood of preferred responses from decreasing. Experiments show that mDPO consistently enhances multimodal LLM performance and reduces hallucination across different model sizes on three widely used benchmarks.

### v2

@@ 1 Introduction
Direct preference optimization (DPO) has emerged as the predominating method for aligning large language models (LLMs) with human preferences Rafailov et al. (2023); Zhao et al. (2024). Building on its success in the language modality, recent studies have extended DPO to multimodal scenarios Li et al. (2023); Yu et al. (2024a); Zhou et al. (2024b); Zhao et al. (2023). However, transferring this approach across modalities presents significant challenges. Merely substituting textual preference data with multimodal preference data does not consistently yield positive outcomes and can exacerbate issues such as hallucinations Li et al. (2023); Sarkar et al. (2024).
@@
While recent efforts in multimodal preference learning focus on improving performance through enhanced multimodal preference data Li et al. (2023); Zhao et al. (2023); Xiao et al. (2024); Zhou et al. (2024b); Pi et al. (2024); Sarkar et al. (2024); Yu et al. (2024b); Deng et al. (2024), we investigate the pitfalls of multimodal DPO from a different perspective. Through controlled comparisons, we discover that multimodal LLMs can achieve similar performance even when all images are removed from the multimodal preference data during DPO (see Figure 1). This counterintuitive finding suggests that the failure of DPO in multimodal scenarios may not be solely attributed to data quality. We attribute this to a systematic gap between the theoretical expectations and practical implementations of the DPO objective in multimodal settings (refer to Figure 2). While DPO aims to compute implicit rewards conditioned on all input modalities, it may prioritize language-only preferences and overlook the image condition (i.e., unconditional preference), leading to suboptimal model performance and increased hallucination. After applying DPO, the model may show an increased tendency to ignore the provided image and generate responses based solely on the question (illustrated in Figure 3).
@@ 3.1 Conditional Preference Optimization
$\mathcal{L}_{\text{CoPO}}=-\log\sigma\Big{(}\beta\log\frac{\pi_{\theta}(y_{w}|m_{w},q)}{\pi_{\text{ref}}(y_{w}|m_{w},q)}-\beta\log\frac{\pi_{\theta}(y_{w}|m_{l},q)}{\pi_{\text{ref}}(y_{w}|m_{l},q)}\Big{)}.$
@@ 4.1 Experimental Setup
Models. We apply mDPO on two multimodal LLMs in different sizes. Bunny-v1.0-3B He et al. (2024) is a 3B model building upon SigLIP Zhai et al. (2023) and Phi-2 Javaheripi et al. (2023). It is pretrained on 2M image-text pairs and finetuned on 695K instruction tuning data. LLAVA-v1.5-7B Liu et al. (2024a) is a 7B model based on CLIP Radford et al. (2021) and Vincuna Chiang et al. (2023). It is pretrained on 558K image-text pairs and finetuned on 665K instruction tuning data.
Preference Data. We sample 10K preference data from Silkie Li et al. (2023) with instructions from LLaVA-Instruct-150K Liu et al. (2024a) for training. The original Silkie dataset contains 80K preference data collected on 12 multimodal LLMs. While the original Silkie paper explores the effect of extreme data size, we follow the majority of prior works using around 10K data for preference optimization Sun et al. (2023); Zhao et al. (2023).
@@
Baselines. We primarily compare mDPO with standard DPO. The standard DPO baseline shares the same training process, data, and hyper-parameters, despite different learning objectives. We further provide the results of other multimodal LLMs for reference, although they are not directly comparable due to different base models, preference data, and alignment methods. This group contains GPT-4V Achiam et al. (2023), LLaVA-v1.5-13B Liu et al. (2024a), Qwen-VL-Chat Bai et al. (2023), POVID Zhou et al. (2024b), HACL Jiang et al. (2024), OPERA Huang et al. (2024), VCD Leng et al. (2024), EOS Yue et al. (2024), HA-DPO Zhao et al. (2023), HALVA Sarkar et al. (2024), RLHF-V Yu et al. (2024a), HSA-DPO Xiao et al. (2024), and Silkie Li et al. (2023). Some of them are contemporary works to ours.
@@ 5 Related Work
Reinforcement learning from human feedback (RLHF; Christiano et al. 2017; Ouyang et al. 2022) has proven to be an effective approach for aligning LLMs with human values. Direct preference optimization (DPO; Rafailov et al. 2023), which involves directly optimizing LLMs based on human preferences, has been widely adopted in RLHF due to its strong performance and the elimination of the need for a separate reward model. Significant efforts have been made to further enhance the efficacy and efficiency of DPO, which can be categorized into algorithmic and data-related advancements. On the algorithmic side, various approaches aim to improve the efficiency of DPO. For example, ORPO Hong et al. (2024) models preferences using an odds ratio and combines instruction fine-tuning and preference optimization into a unified training process. Methods such as CPO Xu et al. (2024a), TPO Saeidi et al. (2024), and SimPO Meng et al. (2024) simplify DPO by eliminating the use of a reference model, thereby reducing computational and memory overhead. Additionally, IPO Azar et al. (2024) addresses the issue of reward overfitting in DPO. On the data side, approaches such as KTO Ethayarajh et al. (2024) and NCA Chen et al. (2024) seek to overcome DPO’s requirement for paired preference data by designing optimization goals that can also utilize unpaired data. Iterative DPO Xu et al. (2023); Yuan et al. (2024); Xiong et al. (2024) and SPPO Wu et al. (2024) propose sampling preference data in an on-policy manner, achieving better results than off-policy DPO. WPO Zhou et al. (2024a) adapts off-policy data to resemble on-policy data by reweighting preference pairs. More closely related to our work, studies by Park et al. (2024) and Dong et al. (2024) address the reward hacking problem in textual preference optimization, where human preference may be biased towards longer outputs. They propose to calibrate the rewards in DPO with respect to output length. In this work, we discover a novel challenge in multimodal DPO, where preference optimization often neglects images. We then propose a solution to this problem through conditional preference optimization.
In multimodal scenarios, recent works mainly focus on creating multimodal preference data Li et al. (2023); Zhao et al. (2023); Xiao et al. (2024); Zhou et al. (2024b); Pi et al. (2024); Sarkar et al. (2024); Yu et al. (2024b); Deng et al. (2024). These efforts include collecting human preference Sun et al. (2023); Yu et al. (2024a), preference from advanced multimodal LLMs Li et al. (2023); Yu et al. (2024b), and preference from the model to align itself Deng et al. (2024). In terms of learning objectives, recent works mainly follows DPO for LLMs Li et al. (2023); Zhao et al. (2023); Zhou et al. (2024b). Some also apply reinforcement learning Sun et al. (2023); Jing and Du (2024) and contrastive learning Sarkar et al. (2024); Jiang et al. (2024). Our work studies an overlooked but crucial problem in the multimodal DPO objective.
@@ 6 Conclusion
We propose mDPO, a preference optimization method dedicated to multimodal scenarios. mDPO leverages conditional preference optimization to encourage multimodal LLMs to capture preference labels based on both visual and language cues. It further introduces anchored preference optimization to prevent the likelihood of preferred responses from decreasing. Experiments show that mDPO consistently enhances multimodal LLM performance and reduces hallucination across different model sizes on three widely used benchmarks. Our method could be extended to reduce hallucinations and enhance trustworthiness across broader multimodal data and scenarios, including multiple images (Wang et al., 2024), videos (Li et al., 2024), in-context learning (Xu et al., 2024b), and risk-sensitive domains (Chaves et al., 2024).
Acknowledgement
We thank the anonymous reviewers for their valuable comments. Fei Wang was supported by the Amazon ML Fellowship. Muhao Chen was supported by the DARPA FoundSci Grant HR00112490370, the NSF of the United States Grant ITE 2333736, and an Amazon Research Award.